Consider the probability distribution of the weekly demand for copier paper (in hundreds of reams) used in a corporation's duplicating center, as shown in the file P04_27.xlsx.

a. Use simulation to generate 500 values of this random variable. Calculate the mean and standard deviation of the given distribution. (Round your answers to two decimal places.)

Mean
Standard Deviation

b. Find the mean and standard deviation of the simulated values. (Round your answers to two decimal places.)

Mean
Standard Deviation

c. Use your simulated values to estimate the probability that weekly copier paper demand is within one standard deviation of the mean. (Round your answers up to one decimal place.)

Mean - Standard Deviation    
Mean + Standard Deviation    
P(demand within 1 Std. Dev. of the Mean) %

Why is this only an estimate, not an exact value?

Here is the info in the excel sheet

Weekly copier paper demand
Demand Probability
10 0.05
11 0.10
12 0.14
13 0.16
14 0.21
15 0.13
16 0.11
17 0.04
18 0.03
19 0.02
20 0.01
Answer #1

Demand Probability Cumulative Probability 10 0.05 12 13 14 15 16 17 18 19 20 0.14 0.16 0.21 0.13 0.11 0.04 0.03 0.02 0.01 0.0

I am unable to show all the 500 simulations here, so I am showing you the first 100 only.

Random # Demand
0.2853049 12
0.1314233 11
0.7103724 15
0.7141029 15
0.0424316 10
0.0432327 10
0.6624186 15
0.6974529 15
0.2928261 13
0.5127866 14
0.7191688 15
0.8743126 16
0.5971292 14
0.8753781 16
0.7053927 15
0.2994464 13
0.1472498 11
0.3052605 13
0.7218783 15
0.9778462 19
0.9357838 17
0.6586554 14
0.532113 14
0.2627841 12
0.3621545 13
0.8058528 16
0.6430508 14
0.4532812 14
0.5949728 14
0.9100599 17
0.6525201 14
0.6520141 14
0.7329046 15
0.6616399 15
0.0692224 11
0.3073655 13
0.9700517 19
0.0719604 11
0.1973078 12
0.1231753 11
0.26169 12
0.8621058 16
0.449521 13
0.133875 11
0.3841196 13
0.6939706 15
0.8619227 16
0.6258181 14
0.7355935 15
0.6808148 15
0.3317924 13
0.5110906 14
0.9698784 18
0.569545 14
0.37036 13
0.9515356 18
0.1448155 11
0.3119094 13
0.3153497 13
0.320733 13
0.1451369 11
0.7896199 15
0.5518004 14
0.3953913 13
0.4387107 13
0.3048158 13
0.4857281 14
0.9524587 18
0.8009903 16
0.620837 14
0.3953379 13
0.4910639 14
0.4067192 13
0.5164161 14
0.0533258 11
0.8263863 16
0.0622823 11
0.7067224 15
0.8678687 16
0.6119928 14
0.8139332 16
0.3917536 13
0.9032413 17
0.1905865 12
0.4529227 14
0.4640331 14
0.4241396 13
0.8264788 16
0.718925 15
0.0985041 11
0.0372518 10
0.8009729 16
0.1876476 12
0.3021737 13
0.5910082 14
0.6032203 14
0.2272108 12
0.8265451 16
0.548562 14
0.5501075 14

Mean, μ = 13.85 (b)        

Standard deviation, σ = 2.16      

(c)        

μ - σ = 11.69       

μ + σ = 16.02       

Number of simulations where the demand lies in the range [11.69, 16.02] = 372

P(11.69 < Demand < 16.02) = 0.7     

(d) This is only an estimate and not an exact value because this is based on just 500 simulations.

[Please give me a Thumbs Up if you are satisfied with my answer. If you are not, please comment on it, so I can edit the answer. Thanks.]

Formula =IF(A43<C$30,10,IF(A43<C$31,11,IF(A43<C$32,12,IF(A43<C$33,13,IF(A43<C$34,14,IF(A43<C$35,15,IF(A43<C$36,16,IF(A43<C$37,17,IF(A43<C$38,18,IF(A43<C$39,19,20))))))))))